find all solutions of the equation algebraically.

36t^4 + 29t^2 - 7 = 0


Answer by stanbon(75887)

find all solutions of the equation algebraically.
36t^4 + 29t^2 - 7 = 0
36t^4 + 36t^2 - 7t^2 - 7 = 0

36t^2(t^2 + 1) -7(t^2+1) = 0
(t^2+1)(36t^2-7) = 0
t^2 = -1 or 36t^2 = 7
t = i or t = -i or sqrt(7)/6 or -sqrt(7)/6
==============================================
cheers,
Stan H.